40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 6 01:09:11 2017, johannes, Update, Computer Scripts / Programs, New set of loss measurements XARM_20170905.pdfYARM_20170905.pdf
    Reply  Mon Sep 11 12:56:40 2017, johannes, Update, Computer Scripts / Programs, lossmap attempts 
       Reply  Mon Sep 11 16:46:09 2017, Steve, Update, PEM, earthquakes M5.4eq.png
          Reply  Tue Sep 19 14:00:41 2017, Steve, Update, PEM, earthquakes 7.1MX5.3J.png
             Reply  Fri Oct 27 09:34:20 2017, Steve, Update, PEM, earthquakes recentEQs.png
                Reply  Fri Nov 10 13:46:26 2017, Steve, Update, PEM, M3.1 local earthquake 3.1M_local_EQ.png
                   Reply  Mon Apr 9 08:39:39 2018, Steve, Update, PEM, M5.3 eq Souther CA M5.3_Santa_Cruz_Is.CA.pngafter_M5.3.pngM5.3vac.png
                Reply  Thu Jan 25 08:27:25 2018, Steve, Update, PEM, earthquakes M4_local_eq.pngM4.pngM5.8Ferndale_CA.png
       Reply  Mon Sep 25 17:56:33 2017, johannes, Update, Computer Scripts / Programs, transmitted power during lossmap 
    Reply  Thu Oct 11 13:40:51 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
       Reply  Fri Oct 12 12:29:34 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
          Reply  Fri Oct 12 20:20:29 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
             Reply  Sat Oct 13 20:11:10 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
                Reply  Mon Oct 15 10:32:13 2018, yuki, Update, Computer Scripts / Programs, loss measurements XARM_20181015_1500.pdfYARM_20181015_1500.pdf
                   Reply  Tue Oct 16 00:44:29 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
                   Reply  Fri Nov 2 19:25:16 2018, gautam, Update, Computer Scripts / Programs, loss measurements 
    Reply  Mon Oct 15 12:52:54 2018, yuki, Update, Computer Scripts / Programs, additional comments 
       Reply  Fri Nov 2 16:42:31 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
          Reply  Mon Nov 5 13:52:18 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
             Reply  Tue Nov 6 10:19:26 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
                Reply  Tue Nov 6 15:23:48 2018, gautam, Update, IOO, IMC problematic MCwonky.png
                   Reply  Tue Nov 6 19:02:35 2018, aaron, Update, IOO, IMC problematic 181106_MCTRANS.jpg
                      Reply  Sat Nov 10 17:40:00 2018, aaron, Update, IOO, IMC problematic 
                         Reply  Mon Nov 12 13:53:20 2018, rana, Update, IOO, loss measurement: oscope vs CDS DAQ 
                         Reply  Thu Nov 15 10:21:07 2018, aaron, Update, IOO, IMC problematic 181115_chansDown.pngPD_noise.png
                            Reply  Fri Nov 16 10:53:07 2018, aaron, Update, IOO, IMC problematic Screenshot_from_2018-11-16_19-22-34.png
                               Reply  Sat Nov 17 18:59:01 2018, aaron, Update, IOO, IMC problematic 
                Reply  Wed Nov 7 05:16:16 2018, yuki, Update, Computer Scripts / Programs, arm loss measuremenents 
Message ID: 14248     Entry time: Fri Oct 12 20:20:29 2018     In reply to: 14245     Reply to this: 14251
Author: yuki 
Type: Update 
Category: Computer Scripts / Programs 
Subject: loss measurements 

I ran the script for measuring arm-loss and calculated rough Y-arm round trip loss temporally. The result was 89.6ppm. (The error should be considered later.)

The measurement was done as follows:

  1. install hardware
    1. Put a PD (PDA520) in anti-symmetric beam path to get ASDC signal.
    2. Use a PD (PDA255) in MC2 box to get tranmitted light of IMC. It is used to normalize the beam power.
    3. Connect those 2 PDs to oscilloscope (IP: 192.168.113.25) and insert an internet cable to it.
  2. measure DARK noise
    1. Block beam going into PDs with dampers and turn off the room light.
    2. Run the script "armloss_dcrefl_acdcpd_scope.py" using "DARK" mode.
  3. measure the ASDC power when Y-arm locked and misaligned
    1. Remove dampers and turn off the room light.
    2. Dither-align the interferometer with both arms locked. Freeze outputs when done. (Click C1ASS.adl>!MoreScripts>ON and click C1ASS.adl>!MoreScripts>FreezeOutputs.)
    3. Misalign ETMX + ITMX. (Just click "Misalign" button.)
    4. Further misalign ITMX with the slider. (see previous study: ITMX needs to be misaligned further. Moving the slider by at least +0.2 is plentiful to not have the other beam interfere with the measurement.)
    5. Start the script "armloss_dcrefl_acdcpd_scope.py" using "ETMY" mode, which does the following:
      1. Resume dithering of the YARM.
      2. Check YARM dither error signal rms with CDS. If they're calm enough, proceed. (In the previous study the rms threshold was 0.7. Now "ETM_YAW_L_DEMOD_I" signal was 15 (noisy), then the threshold was set 17.)
      3. Freeze dithering.
      4. Start a new set of averages on the scope, wait T_WAIT (5 seconds).
      5. Read data (= ASDC power and MC2 trans) from scope and save.
      6. Misalign ETMY and wait 5s. (I added a code which switchs LSC mode ON and OFF.)
      7. Read data from scope and save.
      8. Repeat desired amount of times.
  4. calculate the arm loss
    1. Start the script "armloss_AS_calc.py", whose content is follows:
      • requires given parameters: Mode-Matching effeciency, modulation depth, transmissivity. I used the same value as Johannes did last year, which are (huga)
      • reads datafile of beam power at ASDC and MC2 trans, which file is created by "armloss_dcrefl_acdcpd_scope.py".
      • calculates arm loss from the equation (see 12528 and 12854).

Result:

YARM
('AS_DARK =', '0.0019517200000000003') #dark noise at ASDC 
('MC_DARK =', '0.02792') #dark noise at MC2 trans
('AS_LOCKED =', '2.04293') #beam power at ASDC when the cavity was locked 
('MC_LOCKED =', '2.6951620000000003')
('AS_MISALIGNED =', '2.0445439999999997') #beam power at ASDC when the cavity was misaligned
('MC_MISALIGNED =', '2.665312')

\hat{P} = \frac{P_{AS}-P_{AS}^{DARK}}{P_{MC}-P_{MC}^{DARK}} #normalized beam power 

\hat{P}^{LOCKED}=0.765,\ \hat{P}^{MISALIGNED}=0.775,\ \mathcal{L}=89.6\ \mathrm{ppm}

Comments:

  • "ETM_YAW_L_DEMOD_I_OUTPUT" was little noisy even when the arm was locked.
  • The reflected beam power when locked was higher than when misaligned. It seemed strange for me at first. Johannes suggested that it was caused by over-coupling cavity. It is possible when r_{ETMY}>>r1_{ITMY}.
  • My first (wrong) measurement said the arm loss was negative(!). That was caused by lack of enough misalignment of another arm mirrors. If you don't misalign ITMX enough then the beam or scattered light from X-arm would bring bad. The calculated negative loss would be appeared only when \frac{\hat{P}^{LOCKED}}{\hat{P}^{MISALIGNED}} > 1 + T_{ITM}
  • Error should be considered.
  • Parameters given this time should be measured again. 
ELOG V3.1.3-